Shift report 20020114 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Jan 14 2002 - 15:35:42 EST

  • Next message: Jens Jørgen Gaardhøje: "Draft Agenda for Brahms collaboration meting Jan 28"

    ------------------------------------------------------------
    Date: Mon Jan 14 15:35:42 2002, Shift: 08:00-16:00
    Supervisor: Eun-Joo Kim, with: 
    Log book#12 pages 58 - 59
    
    Shift Summary:
    ------------------------------------------------------------
    1. Previous beam dumped around 10:20am.
       Beam was cogged at 12:45, but quality was not so good.
       Rate was not stable...
       beam for 12:30 - 14:12, about 2hours.
       At 15:15 "Preparing for the next ramp".
    
    2. Pedestal and DVM : Run 6458, 6474, 6476
     
       Data Run : 6457 at FFS@20, MRS@90
                  -> program FFS@20, 1/15A : done
       
       FFS/BFS : moved to 3-deg, 1/3A field setting.
    
       short data run : 6467, 6468, 6473  
       Data run : 6475 (trigger 6 scale down 2)
    
    3. No new detector problem.
       T3 tripped once, and DAQ was not working well.
       During the camac pedestal run, trigger 1 was 
       increased instead of trigger 7. SBE(1-6)bins were Reset.
     
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Mon Jan 14 2002 - 15:36:18 EST